I am having a hard time with this question.
Use the geometric sequence of numbers 1, 2, 4, 8,…to find the following:
a) What is r, the ratio between 2 consecutive terms?
b) Using the formula for the nth term of a geometric sequence, what is the 24th term?
c) Using the formula for the sum of a geometric series, what is the sum of the first 10 terms?
Thanks for your time.

Answer by stanbon(75887)

Use the geometric sequence of numbers 1, 2, 4, 8,…to find the following:
a) What is r, the ratio between 2 consecutive terms?
r=2/1=2
b) Using the formula for the nth term of a geometric sequence, what is the 24th term?
a(n)= ar^(n-1)
a(24)=1*2^23=8,388,608
c) Using the formula for the sum of a geometric series, what is the sum of the first 10 terms?
S(n)=a[(r^(n+1)-1)/(r-1)[
S(24) = 1[(2^11 -1)/(2-1)]
S(24) = 2047
Cheers,
Stan H.